OpenStudy (anonymous):

What is the simplified form of x squared minus 36 all over x squared minus 12 times x plus 36.? x minus 6 all over x plus 6. x minus 6 all over x minus 2. x plus 6 all over x minus 6. x minus 4 all over x plus 2.

OpenStudy (cggurumanjunath):

(X^2-36)/(X^2-12X+36)=(X+6)(X-6)/(X-6)(X-6)=(X+6)/(X-6) IS THE ANSWER

OpenStudy (anonymous):

thanks

OpenStudy (cggurumanjunath):

YOU ARE WELCOME DID YOU GET THE CONCEPT ?